Salley and Bob are eloping but Salley is locked in her 3rd floor bedroom. Their only hope of a wedding is that Salley lets down her 3 story long hair and in the hope the weave holds as he climbs up. Or bob can throw her a rope and let her climb down. They attempt teh rope first. The window is 28ft up and Bob releases the rope with an initial velocity of 30ft/s. the release point is 5 feet above the ground. Will the rope reach the window? Explain using the discriminant. h=-16t squared + v subzero t + s
Answer by Alan3354(69443)

h=-16t squared + v subzero t + s
h(t) = -16t^2 + 30t + 5
t for max height = -b/2a = -30/-32
t = 15/16 seconds
h(15/16) = -225/16 + 450/16 + 5
hmax = 305/16 feet = 19.0625 feet
Not even close. She should let her hair down, tie the rope to it, then pull the rope up.